Implement B(x,y,z) = y’z + x’y’z’ + xz using a 2^3x 1 multiplexer.
Answer #1

BCX,Y,Z) = y2 +xyz +xz zato al al 10 Bla, y, z) = En[0,1,5,7) -0-000-- I Bluny/Z) = y2 + xyz +XZ i Mox .. 17 .